What is the next number in the sequence
7, 21, 8, 72, 9

Answer by greenestamps(13200)



It is NEVER possible, in ANY problem like this, to know the RIGHT answer.

ANY next number would make a valid sequence.

And any next number you find by a logical pattern might still be "wrong", because the author of the problem used a different logical pattern.

Since it is never possible to know the right answer, it is never worthwhile to spend much time on any problem like this. I usually spend no more than a minute or so, and if I can't see anything, I find better ways to spend my time and effort.

In this one, I see the 7-8-9 as a separate sequence; then the other numbers are obtained from those:
7*3 = 21
8*9 = 8*3^2 = 72
so 8*3^3 = 8^27 = 216

But there is no way of knowing if this is the "right" answer....